# CauSkelNet: Causal Representation Learning for Human Behaviour Analysis

### v1

@@ Abstract
Constrained by the lack of model interpretability and a deep understanding of human movement in traditional movement recognition machine learning methods, this study introduces a novel representation learning method based on causal inference to better understand human joint dynamics and complex behaviors. We propose a two-stage framework that combines the Peter-Clark (PC) algorithm and Kullback-Leibler (KL) divergence to identify and quantify causal relationships between joints. Our method effectively captures interactions and produces interpretable, robust representations. Experiments on the EmoPain dataset show that our causal GCN outperforms traditional GCNs in accuracy, F1 score, and recall, especially in detecting protective behaviors. The model is also highly invariant to data scale changes, enhancing its reliability in practical applications. Our approach advances human motion analysis and paves the way for more adaptive intelligent healthcare solutions.
@@ I Introduction
In recent decades, machine learning has made significant strides in the field of sports science, particularly in behavioural analysis and performance evaluation [1][2][3]. The integrated application of computer-aided data processing, mathematical modelling, and simulation optimisation has accelerated the development of behavioural analysis methods and broadened the scope of sports technique exploration [4][5]. These advancements have enabled in-depth analysis of complex behavioural datasets, which was previously challenging using traditional methods [6].
Human behaviour is a multidimensional, complex phenomenon encompassing physiological and psychological aspects, profoundly influenced by emotions and cognition [7]. Research indicates that interactions between behaviours involve intricate Bayesian causal frameworks, which are further modulated by emotional states [7][8]. This complexity necessitates the adoption of more advanced, interpretable algorithmic approaches to comprehensively understand and simulate these relationships.
Despite substantial progress in movement recognition and prediction through machine learning, contemporary research often overlooks rigorous exploration of the underlying causal relationships behind behavioural dynamics. This limitation leads to insufficient model interpretability, thereby impacting our in-depth understanding of human behaviour. Such oversight not only restricts the explanatory power of movement recognition models but also diminishes their potential in guiding optimisation and intervention strategies [6]. For instance, in the field of gait analysis, although machine learning has been widely applied, most studies have yet to delve into the fundamental causal relationships embedded within biomechanical data [9].
To overcome the limitations of existing research, this work proposes an innovative approach combining KL divergence with causal inference, aimed at simulating the complex interactions between human joints in behavioural contexts, thereby learning more effective representations of human kinematic data. We integrate this method with graph convolutional neural networks, determining the efficacy of learned representations through comparison with baseline models. This approach not only pioneers a new application of causal inference learning in the field of human kinematics but also provides a mathematically rigorous framework for elucidating the dynamic interrelationships between joints. This novel representation learning method can reveal joint interactions across different behavioural patterns, deepening our understanding of overall movement patterns and paving the way for new research directions in human-centred computation and behavioural analysis. The core contributions of this research are as follows:
  • Innovative Causal Representation Learning Framework: The study introduces a novel framework that integrates causal inference into human motion analysis. This framework successfully captures the causal relationships between joints, providing a new perspective on representing human movement.
  • Combination of PC Algorithm and KL Divergence: The research combines the PC algorithm with KL divergence in a two-stage causal representation learning method. The PC algorithm is used to uncover potential causal structures, while KL divergence quantifies the strength of mutual causal relationships, resulting in more precise and meaningful motion representations.
@@
  • Highly Interpretable and Robust Motion Recognition: The study develops a motion recognition representation learning method that is both highly interpretable and robust. This approach generates easily comprehensible causal relationship graphs and demonstrates strong invariance to data scale changes, enhancing the model’s reliability and applicability in practical scenarios.
@@ II-A Traditional Kinesiology Research
Traditional kinematic studies primarily employ motion capture and bio-mechanical modelling to analyse human movement mechanisms. These methods quantify the dynamic changes of various body parts during movement, revealing the complex interactions between physical activities and physiological functions [10][11]. However, these approaches have inherent limitations in representation learning. They mainly quantify observable physical indicators (such as joint angles and force output) but fail to effectively learn and represent complex inter-joint relationships. For instance, James et al. [10] demonstrated that fatigue significantly altered bio-mechanical and neuromuscular responses during landing, highlighting the inadequacies of traditional bio-mechanical frameworks in capturing and representing factors of performance variation. Furthermore, Cop et al.[12] found that while the anatomical and functional characteristics of musculoskeletal structures are well understood, the causal interactions between neural and bio-mechanical levels in specific movements remain difficult to represent effectively. These complex physiological and neuromechanical dynamics suggest that traditional methods struggle to comprehensively learn and represent the multidimensional factors influencing motor performance. This underscores the necessity of introducing more advanced representation learning methods in kinematic studies to better capture and understand the intrinsic mechanisms of human movement.
@@ II-B Multi-modal Data and Representation Learning
Advancements in deep learning and multi-modal data modelling have significantly enhanced our understanding of human behaviour. Researchers have developed more sophisticated and effective behavioural representation methods by utilising various data sources, including motion capture, bio-mechanical analysis, and physiological monitoring. The implementation of Graph Convolution Networks (GCN) and Parallel Spatiotemporal Encoding Motion Representation (P-STEMR) models has paved new avenues for analysing movement dynamics [2]. GCN optimises joint data processing, effectively learning and representing inter-joint interactions, and exploring potential pathways influencing movement dynamics. Concurrently, the Human Activity Recognition-Protective Behaviour Detection (HAR-PBD) framework has enhanced the ability to represent complex behavioural patterns through continuous monitoring of joint dynamics and interactions [13]. However, existing research primarily focuses on unidirectional influences, overlooking the complex interplay and causal relationships between behaviour, emotions, and other factors. This limitation underscores the need for more in-depth causal analysis to develop advanced learning models capable of comprehensively representing the complexity of human behaviour. Such an approach can not only improve the accuracy of behaviour recognition but also provide a richer informational basis for developing personalised intervention strategies.
@@ III-A Formulation
Let $\mathcal{G}=(V,E)$ be a graph representing the human body structure, where $V=\{v_{1},\ldots,v_{n}\}$ denotes joints and $E$ denotes connections between joints. Our objective is to derive a representation $\phi:V\rightarrow\mathbb{R}^{d}$ that encapsulates the causal relationships between joints.
We employ the PC algorithm to elucidate the causal structure. For variables $X,Y\in V$ and a set of conditioning variables $S\subset V\setminus\{X,Y\}$, we examine the conditional independence with Eq.1. The PC algorithm yields a completed partially directed acyclic graph (CPDAG) $\mathcal{G}^{*}=(V,E^{*})$.
To quantify the strength of causal relationships and derive joint representations, we utilize KL divergence. For any edge $(i,j)\in E^{*}$, we define:
@@
where $\text{Ne}(i)$ denotes the set of neighbors of node $i$ in $\mathcal{G}^{*}$.
@@ III-B Theoretical Guarantees
3:Bayesian Network Formation: Formulate DAG $\mathcal{G}=(V,\vec{E})$ by directing $E$ using dependencies derived from $\mathcal{D}$.
@@ IV-A Dataset
The EmoPain dataset [23] employed in this study is designed to analyze movement behaviors and pain recognition in chronic pain (CP) patients relative to a healthy control group [24]. This dataset records pain levels and protective behaviors from 18 IMUs of 12 healthy and 18 CP participants within a controlled movement challenge, providing insights into how pain influences movement. As depicted in Figure 1, the dataset constructs a fundamental human causal graph using the X, Y, and Z coordinates of 22 body joints, capturing body postures during movements and facilitating in-depth analyses of human motion patterns.
@@ IV-B Experiment Design
In this study, depicted in Figure 1, we used the EmoPain dataset to create an undirected causal graph model of human structure, examining causal connections between joints in protective and non-protective behaviors and their impact on pain analysis. We processed the three-dimensional joint coordinate data (X, Y, Z) to capture dynamic interactions, then applied the PC algorithm to learn a DAG of potential causal relationships. A Bayesian network was constructed, and conditional probability distributions were determined using maximum likelihood estimation. For each joint pair (e.g., Joint 1 and Joint 2), we calculated the conditional probability distribution and used KL divergence to measure information transfer efficiency, reflecting the reduction in uncertainty of one joint’s state given another’s. We compared KL divergence across dimensions to identify the dominant causal direction, selecting the most frequent direction and using its average KL divergence as a weight for information transfer efficiency. This analysis delved into joint interactions and their role in pain perception and protective behaviors, offering new insights for behavior-based pain analysis.
@@ IV-C Experiment Analysis
Figure 3: Causal graphs for four different exercise types in both protective and non-protective behaviors are displayed, from left to right: One leg stand, Reach forward, Bend, and Others. Each colored box highlights the differences between protective and non-protective behaviors based on graph representations. For instance, the direction of the arrows are different.
@@
TABLE I: Performance Metrics for Baseline vs. Causal GCN in test
Exercise Type Baseline Causal GCN
Acc. Prec. F1 Rec. Acc. Prec. F1 Rec.
One Leg Stand 0.933 0.895 0.862 0.910 0.942 0.903 0.886 0.969
Bend 0.883 0.894 0.859 0.792 0.916 0.915 0.908 0.920
Reach Forward 0.882 0.899 0.826 0.736 0.892 0.897 0.849 0.794
Other 0.939 0.924 0.474 0.320 0.940 0.878 0.569 0.439
@@ V-B Evaluation and Result
As shown in Table I, Causal GCN outperformed baseline models across various movement types. For single-leg standing, Causal GCN achieved superior F1 score (0.886) and recall (0.969) compared to the baseline (0.862 and 0.910 respectively). In ”bending” activities, it improved recall from 0.792 to 0.920. For complex ”other” movements, F1 score increased from 0.474 to 0.569 and recall from 0.320 to 0.439.
Figure 4 illustrates Causal GCN’s consistent performance advantage across all movement types. This superiority in accuracy, F1 score, and recall demonstrates its effectiveness in capturing subtle joint dynamics and behavioral relationships.
These results highlight the benefits of integrating causal structures into behavioral analysis models, enhancing both interpretability and performance. The Causal GCN’s adaptability to different movement types and its ability to capture behavior changes associated with pain make it promising for developing personalized treatment strategies.
@@
TABLE II: Performance Metrics Comparison of General Models without Exercise Type Labels
Model Accuracy Precision F1 Recall
Baseline - Test 0.928 0.900 0.724 0.625
Causal GCN - Test 0.942 0.908 0.793 0.733
To assess the generalization capability of causal graph models in behavior recognition, we compared the performance of the baseline model and the general causal graph model without exercise type labels. As shown in Table II, the Causal GCN model outperformed the baseline across all metrics over 50 epochs of evaluation. Causal GCN significantly improved performance metrics compared to the baseline model: testing F1 score (0.793 vs. 0.724), and testing recall (0.733 vs. 0.625). Figure 6 illustrates the Causal GCN’s superior and more stable performance across all indicators. These results demonstrate the model’s enhanced capability in identifying protective behaviors and its adaptability across different behavior types, confirming that integrating causal structures improves both generalizability and interpretability in behavior recognition tasks.
VI Discussion
This study introduces an innovative causal representation learning approach, advancing the field of human motion and behavior analysis. By synergizing causal inference with state-of-the-art representation learning techniques, we propose a novel dimensional perspective for information extraction, thereby enhancing model performance, interpretability, and generalization capabilities. This methodology demonstrates significant efficacy in feature extraction from complex human data, substantially improving the performance and interpretability of human-centered computational models. Our approach underscores the critical role of causal reasoning in representation learning, emphasizing that elucidating deeper feature relationships beyond mere statistical correlations is paramount for comprehending intricate human behavior patterns[25]. The findings suggest that capturing causal structures in data can lead to more robust and generalizable models, particularly in the context of human behavior analysis. Future research directions may include the integration of more sophisticated causal discovery algorithms into the representation learning framework, potentially revolutionizing the analysis of complex human data across various domains, including healthcare, human-computer interaction, and affective computing.
Furthermore, the proposed causal representation learning method shows potential advantages in analyzing the differential causal relationships between protective and non-protective joint behaviors, with findings aligning with existing medical literature on the impact of pain and emotional distress on behavior. Through the analysis of the EmoPain dataset, this study identified behavior patterns consistent with several established medical models: in line with Vlaeyen and Linton’s fear-avoidance model, chronic low back pain patients exhibited stronger causal relationships between hip and knee joints during bending tasks[26]; similarly, consistent with Sullivan et al.’s pain catastrophizing theory, participants with high pain catastrophizing showed increased causal strength between trunk and lower limb joints during sit-to-stand tasks[27]. Additionally, supporting Hodges and Tucker’s motor adaptation model, chronic pain patients demonstrated more complex causal networks among spinal joints during extension movements[28]. This method also provides a novel perspective for identifying pain-related fear-avoidance behaviors as described by Asmundson et al. , and by integrating emotional score data, it captures the interaction between emotional states and chronic pain emphasized by Geisser et al., such as the generalized reduction in causal joint strength observed in participants with high depression levels[29][30]. These findings not only offer a new data-driven approach for validating existing medical theories but also provide potential applications for quantifying emotion-behavior relationships and developing personalized pain management strategies.
VII Conclusion

### v3

@@ Abstract
Traditional machine learning methods for movement recognition often struggle with limited model interpretability and a lack of insight into human movement dynamics. This study introduces a novel representation learning framework based on causal inference to address these challenges. Our two-stage approach combines the Peter-Clark (PC) algorithm and Kullback-Leibler (KL) divergence to identify and quantify causal relationships between human joints. By capturing joint interactions, the proposed causal Graph Convolutional Network (GCN) produces interpretable and robust representations. Experimental results on the EmoPain dataset demonstrate that the causal GCN outperforms traditional GCNs in accuracy, F1 score, and recall, particularly in detecting protective behaviors. This work contributes to advancing human motion analysis and lays a foundation for adaptive and intelligent healthcare solutions.
@@ I Introduction
Over the decades, traditional kinematic studies, using techniques like motion capture and biomechanical modeling, have advanced our understanding of human movement [1][2]. While valuable for analyzing body segment dynamics in areas like rehabilitation and performance evaluation, these methods often rely on observable metrics, such as joint angles and forces, limiting their ability to capture complex inter-joint dynamics. Recently, machine learning has transformed sports science by enabling advanced analysis of complex behavioral datasets [3][4]. Techniques like deep learning and reinforcement learning have improved behavioral analysis and expanded exploration of sports techniques, addressing challenges such as modeling multidimensional joint interactions and uncovering their causal relationships.
Human behavior is a complex interplay of physiological and psychological dimensions shaped by emotions and cognition [5]. Traditional frameworks offer foundational insights but often neglect the causal, bidirectional relationships between behavioral components. Research indicates that behavioral interactions follow Bayesian causal frameworks influenced by emotional and physiological states [5][6]. This underscores the need for advanced, interpretable algorithms to reveal the causal mechanisms driving behavioral dynamics, enabling a deeper understanding of physical and emotional interactions in real-world contexts.
While machine learning excels in movement recognition and prediction, many approaches prioritize classification accuracy over the causal relationships underlying behavioral dynamics [7]. For instance, gait analysis often focuses on biomechanical classification, overlooking bidirectional joint interactions critical for movement optimization [8]. Understanding these causal links is key to improving model interpretability and enabling targeted interventions, paving the way for more personalized and adaptive applications in sports and healthcare.
This study introduces a novel approach combining KL divergence with causal inference to quantify causal relationships between human joints in behavioral contexts. While graph convolutional networks (GCNs) effectively represent inter-joint dynamics, their application in causal modeling remains underexplored. By integrating GCNs with causal inference, this research provides a robust framework for analyzing joint interactions, addressing existing limitations, and enabling advancements in personalized healthcare and sports performance optimization. The key contributions are summarized as follows:
  • Causal Representation Framework: A framework incorporating causal inference into human motion analysis to capture joint relationships and provide new insights into movement.
  • PC Algorithm and KL Divergence Integration: A two-stage causal learning method combining the PC algorithm and KL divergence to uncover causal structures and quantify joint interactions, leading to more precise motion representations.
@@
  • Interpretable and Robust Motion Recognition: A motion recognition approach that generates interpretable causal graphs and demonstrates strong invariance to data scale changes, enhancing reliability and applicability in practical scenarios.
@@ II-A Traditional Kinesiology Research
Traditional kinematic studies have advanced understanding of human movement using techniques like motion capture and biomechanical modeling to quantify dynamic changes in body segments. These methods have contributed significantly to fields such as clinical rehabilitation and sports performance evaluation [2][1], revealing interactions between physical activities and physiological functions. However, they often focus on observable metrics, such as joint angles and force outputs, limiting their ability to model complex inter-joint dynamics. For instance, James et al. [2] showed that fatigue alters biomechanical and neuromuscular responses during landing, while Cop et al. [9] highlighted the difficulty of representing causal interactions between neural and biomechanical levels. While traditional methods provide a strong foundation, they struggle to capture the multidimensional factors influencing motor performance, underscoring the need for advanced representation learning techniques to model these complex dynamics more effectively.
@@ II-B Multi-modal Data and Representation Learning
Advancements in deep learning and multi-modal data modeling have greatly improved human behavior analysis by leveraging diverse data sources such as motion capture, biomechanical analysis, and physiological monitoring. Techniques like Graph Convolution Networks (GCNs) and Parallel Spatiotemporal Encoding Motion Representation (P-STEMR) have significantly advanced the modeling of movement dynamics [3, 10, 11]. GCNs excel in capturing inter-joint interactions, while frameworks like HAR-PBD enhance behavior recognition by continuously monitoring joint dynamics [10]. Despite these successes, most research emphasizes unidirectional influences, neglecting the complex causal relationships between behavior, emotion, and physiology [12, 13]. Integrating causal inference into these models could address this gap, improving the understanding of movement mechanisms. Hybrid approaches, such as combining GCNs with temporal models like LSTMs, have shown promise in capturing dynamic, multimodal relationships [11, 14]. These advancements underscore the need for models that effectively represent behavioral complexity while enabling personalized intervention strategies.
@@ III Methodology
where $X$ and $Y$ represent the state variables (e.g., position, velocity, or acceleration) of two joints in our model, and $S$ is a conditioning set of other joint states that might influence the relationship between $X$ and $Y$. This conditional independence test checks if $X$ and $Y$ are independent given $S$. If $X\perp\!\!\!\perp Y\mid S$ holds, it indicates no direct causal link between $X$ and $Y$ after accounting for the influence of $S$. This test is key to constructing the initial undirected graph in our causal model. For example, when analyzing the relationship between the knee joint ($X$) and ankle joint ($Y$), conditioning on the hip joint state ($S$) helps determine if their relationship is direct or mediated by the hip joint.
@@
where $D_{\text{KL}}(P\parallel Q)$ measures the divergence between two distributions, $P(x)$ and $Q(x)$, representing the probability of a joint’s state under conditions $P$ and $Q$, respectively. A larger $D_{\text{KL}}$ indicates a greater difference between the distributions, suggesting a stronger potential causal relationship. In our model, this is used to assess the asymmetric information flow between joints. For example, if $D_{\text{KL}}(P_{i}\parallel P_{j})$ is significantly larger than $D_{\text{KL}}(P_{j}\parallel P_{i})$, it suggests joint $i$’s movement pattern provides more information about joint $j$’s movement, implying a potential causal direction from $i$ to $j$.
@@ III-A Formulation
Let $\mathcal{G}=(V,E)$ be a graph representing the human body structure, where:
  • $V=\{v_{1},\ldots,v_{n}\}$ denotes the set of nodes representing human joints, with $n$ being the total number of joints;
  • $E\subseteq V\times V$ denotes the set of edges representing physical or causal connections between joints;
  • $v_{i}\in V$ represents the $i$-th joint in our skeletal model.
Our objective is to derive a representation $\phi:V\rightarrow\mathbb{R}^{d}$ that encapsulates the causal relationships between joints, where $d$ is the dimension of our representation space. We employ the PC algorithm to elucidate the causal structure. For variables $X,Y\in V$ and a set of conditioning variables $S\subset V\setminus\{X,Y\}$, we examine the conditional independence with Eq.1. The PC algorithm yields a completed partially directed acyclic graph (CPDAG) $\mathcal{G}^{*}=(V,E^{*})$, where $E^{*}$ represents the set of directed edges identifying causal relationships. To quantify the strength of causal relationships and derive joint representations, we utilize KL divergence. For any edge $(i,j)\in E^{*}$, we define:
@@
where:
  • $x_{i},x_{j}$ represent the state variables of joints $i$ and $j$ respectively;
  • $\mathcal{X}_{j}$ is the set of all possible states for joint $j$;
  • $P(x_{j}\mid x_{i})$ is the conditional probability of joint $j$ being in state $x_{j}$ given joint $i$ is in state $x_{i}$;
  • $P(x_{j})$ is the marginal probability of joint $j$ being in state $x_{j}$.
@@
where:
  • $\text{Ne}(i)$ denotes the set of neighboring joints directly connected to joint $i$ in $\mathcal{G}^{*}$;
  • The representation $\phi(v_{i})$ is a vector whose components are normalized KL divergence values;
  • Each component represents the relative strength of causal influence from joint $i$ to its neighboring joints.
@@ III-B Theoretical Guarantees
3:DAG Formation: Formulate DAG $\mathcal{G}=(V,\vec{E})$ by directing $E$ using dependencies derived from $\mathcal{D}$.
@@ IV-A Dataset
The EmoPain dataset [24] employed in this study is designed to analyze movement behaviors and pain recognition in chronic pain (CP) patients relative to a healthy control group [25]. This dataset records pain levels and protective behaviors from 18 IMUs of 12 healthy and 18 CP participants within a controlled movement challenge, providing insights into how pain influences movement. As depicted in Figure 1, the dataset constructs a fundamental human causal graph using the X, Y, and Z coordinates of 22 body joints, capturing body postures during movements and facilitating in-depth analyses of human motion patterns. EmoPain, as a publicly available dataset, provides detailed demographic information about the participants in its referenced articles[24]. Specific details can also be found in Table I in the Appendix.
@@ IV-B Experiment Design
In this study, depicted in Figure 1, we used the EmoPain dataset to create an undirected causal graph model of human structure, examining causal connections between joints in protective and non-protective behaviors and their impact on pain analysis. We processed the three-dimensional joint coordinate data (X, Y, Z) to capture dynamic interactions, then applied the PC algorithm to learn a DAG of potential causal relationships. A DAG was constructed based on the PC algorithm output, and the strength of causal relationships between joints was quantified using KL divergence measures. For each joint pair (e.g., Joint 1 and Joint 2), we calculated the conditional probability distribution and used KL divergence to measure information transfer efficiency, reflecting the reduction in uncertainty of one joint’s state given another’s. We compared KL divergence across dimensions to identify the dominant causal direction, selecting the most frequent direction and using its average KL divergence as a weight for information transfer efficiency. This analysis delved into joint interactions and their role in pain perception and protective behaviors, offering new insights for behavior-based pain analysis.
@@ IV-C Experiment Analysis
Figure 3: Causal graphs for four representative exercise types in both protective and non-protective behaviors are displayed, from left to right: One Leg Stand, Reach Forward, Bend, and Others. Each colored box highlights the differences between protective and non-protective behaviors based on graph representations, such as the direction and strength of the arrows. Detailed visualizations of all causal graphs for the nine movement types can be found in Appendix C.
@@ V-B Evaluation and Result
As shown in Table LABEL:tab:metrics_comparison, the Causal Graph Convolutional Network (Causal GCN) demonstrates significant effectiveness compared to baseline models across various movement types. The performance metrics—accuracy, macro precision, F1 score, and recall—provide a comprehensive assessment of the model’s capabilities during both the training and testing phases, with activities such as single-leg standing, bending, and reaching forward. In the single-leg standing task, the Causal GCN achieved high performance during the testing phase, with an F1 score of 0.883 and a recall of 0.9687, the model not only exhibits high precision in classifying the correct instances but also demonstrates a strong ability to retrieve all relevant instances. These results significantly surpass those of the baseline model, which reported an F1 score of 0.8578 and recall of 0.912. This improvement indicates that the Causal GCN effectively captures the behavioral causal relationships triggered by pain-induced protective actions. By accurately modeling the subtle joint dynamics involved in protective behaviors, the Causal GCN can better distinguish between normal and pain-affected movements.
In the bending task, the Causal GCN improved the recognition of positive samples. The recall in the training phase increased from 0.7731 in the baseline model to 0.9328 in the Causal GCN. This improvement indicates a greater ability to identify actual instances of the target behavior, thus reducing the number of false negatives. Similarly, in the complex ”other” movement category, which involves a large and diverse dataset, the Causal GCN outperformed the baseline model. The training phase F1 score increased from 0.4738 to 0.5758, and recall rose from 0.3214 to 0.447. These results suggest that the model is more capable of handling complex and variable real-world data, making it more reliable for practical applications.
Across all categories, the Causal GCN consistently consistently outperformed the baseline in accuracy during both training and testing phases. This stability and high accuracy underscore its potential as a reliable model for motion behavior classification and pain detection tasks. The observed improvements in F1 score and recall in the ”other” category indicate its superior ability to interpret complex motion and causal relationships. Unlike traditional models that may struggle with understanding the intricate connections between joints and behaviors, the Causal GCN excels in analyzing causal flows and using this information to make accurate predictions.
To further validate our findings, Figure 4 visually illustrates the performance trends. The Causal GCN’s metrics consistently rank higher and show greater stability across all movement types compared to the baseline model. This graphical representation provides strong evidence of its overall performance advantage, reinforcing the conclusion that integrating causal structures into the model enhances both interpretability and performance.
When compared to other studies in human motion analysis and pain detection, our Causal GCN model stands out. While some previous models focus solely on improving accuracy without considering the underlying causal mechanisms, our approach not only achieves high accuracy but also offers valuable insights into the causal relationships among joints. This ability to model causal connections leads to a deeper understanding of how pain influences human motor behavior, providing a solid foundation for developing personalized pain management and treatment strategies.
@@
To evaluate the generalization capability of causal graph models in behavior recognition, we conducted an additional experiment comparing the performance of the baseline model with that of the general causal graph model, which did not rely on exercise type labels. As shown in Table LABEL:tab:general_model_metrics, after 50 epochs of evaluation, the Causal GCN outperformed the baseline model across all metrics. Specifically, the testing F1 score improved from 0.724 to 0.793, and the testing recall increased from 0.625 to 0.733. Figure 5 further illustrates the Causal GCN’s superior and more stable performance across all indicators. These results clearly demonstrate that incorporating causal structures enhances the model’s ability to generalize across different behavior types, making it more adaptable to a wide range of real-world scenarios.
VI Conclusion